# HERACLES++: a multi-dimensional Eulerian code for exascale computing

### v1

@@ 2.5 Monitoring of internal energy in highly supersonic regions
The evolution of the internal energy is then given by :
@@
The condition for $\alpha$ is:
$\alpha=\left\{\begin{aligned} &1\quad&\textrm{if}\quad e\geq\varepsilon\,e_{k}\\ &\Big{(}\frac{e}{\varepsilon\,e_{k}}\Big{)}^{2}\quad&\textrm{if}\quad\varepsilon\,e_{k}>e>\frac{\varepsilon\,e_{k}}{10}\\ &0\quad&\textrm{if}\quad e\leq\frac{\varepsilon\,e_{k}}{10}\end{aligned}\right.$ (12)
@@ 3.1 Numerical discretization
Our notation for space-time discretization is such that the position at the center of the cell is denoted by $r_{i}$ and the mesh nodes are given by $r_{i\pm 1/2}$. The cell size is $\delta r$. The time interval between $t^{n}$ and $t^{n+1}$ is $\delta t$. The notation $U^{n}_{i}$ represents the average quantity at time $t^{n}$ for the cell centered at $r_{i}$. This notation is trivially extended to all geometries and in several dimensions. When using spherical coordinates, $i$ refers to the radius coordinate, $j$ to the polar-angle coordinate and $k$ to the azimuthal-angle coordinate.
@@ 3.1.1 Mesh
In HERACLES++, the computational domain is defined as a logically Cartesian mesh that can be either in Cartesian or spherical coordinates and in 1-D, 2-D, or 3-D. The surface elements ($dS$) and volumes ($dV$) of cells are computed based on the chosen geometry. The grid spacing in each direction can be either regular or logarithmic. Users also have the option to implement a custom grid spacing. It is also possible to initialize the simulation from an HDF5 file containing both the physical quantities and the grid spacing.
In addition, the code includes a moving grid feature that allows to shift the grid along a specified coordinate axis during the simulation. This functionality allows to track regions of interest defined by the user. In the context of SN ejecta, the grid can be allowed to move in the radial direction in order to follow the expanding ejecta. With this option, the grid size can be reduced or resolution can be enhanced in specific regions. Furthermore, the elimination of cells in the inner ejecta reduces the constraints on the CFL conditions.
@@ 3.1.2 Momentum equation discretization in spherical coordinates
Equations 6 are written in conservative form. However, in spherical coordinates, the tensor operator (Eq. 1b) presents geometrical source terms. We use a classical discretization method that conserves angular momentum. The azimuthal component $\phi$ of the momentum equation is then written as:
$\begin{split}\frac{\partial\rho u_{\phi}}{\delta t}+\frac{1}{r^{2}}\frac{\partial r^{2}\rho u_{\phi}u_{r}}{\partial r}+\frac{1}{r\sin\theta}\frac{\partial\sin\theta\,\rho u_{\phi}u_{\theta}}{\partial\theta}\\ +\frac{1}{r\sin\theta}\frac{\partial(\rho u_{\phi}^{2}+P_{\phi\phi})}{\partial\phi}=-\,\Big{(}\frac{\rho u_{\phi}u_{r}}{r}+\frac{\cot\theta\,\rho u_{\phi}u_{\theta}}{r}\Big{)}\end{split}$ (13)
@@
$\begin{split}\frac{\rho u_{\phi}u_{r}}{r}=\frac{r_{i+1/2}-r_{i-1/2}}{(r_{i+1/2}+r_{i-1/2})\,dV}\Big{(}[\rho u_{\phi}u_{r}]_{i+1/2}\,dS_{i+1/2}^{r}\\ +[\rho u_{\phi}u_{r}]_{i-1/2}\,dS_{i-1/2}^{r}\Big{)}\\ \frac{\cot\theta\,\rho u_{\phi}u_{\theta}}{r}=\frac{\sin\theta_{j+1/2}-\sin\theta_{j-1/2}}{(\sin\theta_{j+1/2}+\sin\theta_{j-1/2})\,dV}\\ \times\Big{(}[\rho u_{\phi}u_{\theta}]_{j+1/2}\,dS_{j+1/2}^{\theta}+[\rho u_{\phi}u_{\theta}]_{j-1/2}\,dS_{j-1/2}^{\theta}\Big{)}\end{split}$ (14)
where $dS^{r}$ is the surface elements in the radial direction, $dS^{\theta}$ in the polar direction and $dV$ the volume element. The values at the cell interfaces are computed using outputs from the Riemann solver used for the cell update.
The corresponding equation for the polar component $\theta$ is:
$\begin{split}\frac{\partial\rho u_{\theta}}{\delta t}+\frac{1}{r^{2}}\frac{\partial r^{2}\rho u_{\theta}u_{r}}{\partial r}+\frac{1}{r\sin\theta}\frac{\partial\sin\theta(\rho u_{\theta}^{2}+P_{\theta\theta})}{\partial\theta}\\ +\frac{1}{r\sin\theta}\frac{\partial\rho u_{\theta}u_{\phi}}{\partial\phi}=-\,\Big{(}\frac{\rho u_{\theta}u_{r}}{r}-\cot\theta\,\frac{\rho u_{\phi}^{2}}{r}\Big{)}\end{split}$ (15)
@@
$\begin{split}\cot\theta\,\frac{\rho u_{\phi}^{2}}{r}=\frac{([\rho u_{\phi}^{2}]_{k+1/2}+[\rho u_{\phi}^{2}]_{k-1/2})\cos[(\theta_{j+1/2}+\theta_{j-1/2})/2]}{2\sin[(\theta_{j+1/2}+\theta_{j-1/2})/2]}\\ \times\frac{(dS_{j+1/2}^{\theta}-dS_{j-1/2}^{\theta})}{2\,dV}\end{split}$ (16)
@@ 3.2 Algorithmic steps
HERACLES++ was designed to easily incorporate a variety of numerical schemes. In this paper we concentrate on Godunov-type approach (Godunov & Bohachevsky, 1959) and more specifically on the Monotone Upstream-centered Schemes for Conservation Laws (MUSCL)-Hancock method (van Leer, 1977a, b, 1979, 1984; Toro, 2009) to solve the hydrodynamical equations. This method is a simple but practical choice for achieving second-order accuracy in numerical simulations. Its ability to accurately represent gradients and discontinuities, its robustness, along with its conservation properties, make it a reliable tool to treat complex flows involving strong shocks.
@@
$\tilde{U}^{n}_{L}=\tilde{U}^{n}_{i}-\frac{\delta r_{i}}{2}\sigma\quad\textrm{and}\quad\tilde{U}^{n}_{R}=\tilde{U}^{n}_{i}+\frac{\delta r_{i}}{2}\sigma$ (17)
where $\tilde{U}_{L}$ and $\tilde{U}_{R}$ are the reconstructed primitive variables ($\rho,\vec{u},P$) at the left and right sides of the cell, and $\sigma$ the slope defined using a slope limiter. Three different slope limiters are implemented in HERACLES++: Minmod (Sandham & Yee, 1989), Van Leer, and Van Albada (Berger & Aftosmis, 2005; Toro, 2009). The user can choose the slope limiter at runtime and it is also possible to remain at first order. From these primitive reconstructed variables, we can obtain the reconstructed conservative variables $U_{L,R}$.
@@
$U^{n+1/2}_{L,R}=U^{n}_{L,R}+\frac{\delta t}{2\delta r_{i}}\Big{[}F(U^{n}_{L})-F(U^{n}_{R})\Big{]}$ (18)
where $F$ is the flux. The fluxes are computed from these space-time extrapolated values using the analytical flux function. In 3-D, the reconstruction is done in a similar manner (e.g., see Section 16.5 of Toro 2009).
@@
$U_{i}^{n+1}=U_{i}^{n}+\frac{\delta t}{\delta r}\Big{(}\mathcal{F}_{i-1/2}^{n+1/2}-\mathcal{F}_{i+1/2}^{n+1/2}\Big{)}$ (19)
@@ 3.3 Implementation
HERACLES++  is designed with an emphasis on both performance and portability. This enables the use of a single source code that can be compiled across various architectures (GPUs, CPUs) while ensuring optimal performance on each. The code is modular, making it easy to implement and switch between different slope limiters and Riemann solvers for the flux calculation.
@@ 3.3.1 Parallelism
The code can leverage MPI (Message Passing Interface) for distributed memory parallelism and the Kokkos library (Trott et al., 2022) for shared memory parallelism. The global computational domain is partitioned into equal subdomains, with each subdomain assigned to a separate MPI process. Within each subdomain, the Kokkos library is used to take advantage of shared memory parallelism.
@@
where U_new, U, and S are all of the Kokkos::View type, which is the fundamental data type used in the Kokkos library. The MDRangePolicy specifies how to iterate over a multi-dimensional container.
@@ 3.3.2 Inputs and outputs
The Input/Output (I/O) operations in HERACLES++ are handled by PDI, [^2] PDI stands for Portable Data Interface. Information can be found at https://pdi.dev/master/ a library designed to decouple high-performance simulation codes from I/O management. The simulation generates a significant amount of data, which is output by PDI into HDF5 files. We can initialize a simulation by directly assigning values to the physical fields or by reading an HDF5 file containing the values for these fields. This approach allows us to restart a simulation from a specific checkpoint by loading the corresponding data file.
@@ 4 Benchmarking tests
We have conducted a large number of tests to validate and benchmark HERACLES++. These include the standard suite of hydrodynamical tests for shocks, advection, implosion, or explosion. In Appendix C, we present results for a subset of such tests, in particular those that were performed in 2-D and 3-D. In this section, we present results for the tests that have a more direct relevance to the context of multi-dimensional simulations of stellar explosions. Hence, we present results for a 2-D Kelvin-Helmholtz instability test in Section 4.1, for a 3-D Rayleigh-Taylor instability test in Section 4.2, and for a 3-D Sedov blast wave test in Section 4.3. We then move on to discuss a more astrophysics-oriented test in Section 5 with the simulations in 1-D and 3-D of the propagation of a SN shock in a RSG star envelope.
Figure 1: Colormap of the density for the 2-D Kelvin-Helmholtz instability test and shown at times 00 (left), $3$ (middle) and $6$ (right). The grid resolution corresponds to $[n_{x},n_{z}]=[2048,1024]$. (See Section 4.1 for discussion.)
@@ 4.1 Two-dimensional Kelvin-Helmholtz instability test
The Kelvin-Helmholtz instability test is useful to investigate the development of turbulence arising from a velocity shear within a fluid. Following Lecoanet et al. (2016), the initial conditions in the $(x,z)$ plane for the density $\rho$, the $x$-velocity $u_{x}$ and the $z$-velocity $u_{z}$ are given by:
@@
with $a=0.05$, $\sigma=0.2$, $z_{1}=0.5$ and $z_{2}=1.5$. The initial pressure $P_{0}$ is 10 and we assume an ideal gas with $\gamma=5/3$. The boundary conditions are periodic in every direction.
Figure 2: Instability growth rate for the 2-D Kelvin-Helmholtz instability test shown in Fig. 1. This growth rate is estimated from the evolution of the quantity $\partial\ln u_{z}^{max}/\partial t$ and for three different grid choices, corresponding to “Low” (blue), “Med” (red), and “High” (green) resolution (i.e., $[n_{x},n_{z}]=[512,256]$, $[1024,512]$, and $[2048,1024]$). The dashed line corresponds to the best fitting exponential to the growth rate in the linear regime, which breaks down at $t\sim 2.6$.
Figure 3: Passive scalar concentration $f_{\textsf{x}}$ for the 3-D Rayleigh-Taylor instability test at a time of $5$ (left) and $10$ (right). The Cartesian grid is characterized by $[n_{x},n_{y},n_{z}]=[512,512,1024]$. (See §4.2 for discussion.)
Figure 1 shows the evolution of the density at three times ($t=0,3,6$). The instability results in the creation of small, alternate vortices, which cascade into turbulence. In Fig. 2, we quantify the growth rate of the instability-generated structures by evaluating the evolution of the quantity $\partial\ln u_{z}^{max}/\partial t$, with $u_{z}^{max}$ the maximum velocity at each time, and for different resolutions: $512\times 256$ (blue), $1024\times 512$ (red) and $2048\times 1024$ (green). The linear phase of the instability ends at $t\sim 2.6$ and the black dashed line represents the best exponential fitting for this phase, corresponding to $0.01e^{1.9t}$. The growth rate flattens after that time and remains roughly constant until the end of the simulation, largely independent of the resolution.
We can compare our results to those published by Skinner et al. (2019). The linear evolution that they obtain ends around a time of 2.5 with a value of 1.5, which are similar to our results. Their growth rate is between 1.5 and 2, whereas ours is a little lower between 1.2 and 1.7. When comparing Fig. 1 with the counterpart in their work, we observe fewer structures and a more diffusive scheme, which can explain the lower values in our growth rate. In both simulations, the resolution affects modestly the development of turbulence. Indeed, we only obtain a slight increase in growth rate when employing a higher resolution.
@@ 4.2 Three-dimensional Rayleigh-Taylor instability test
The first three dimensional test we present is the development of the nonlinear R-T instability. The initial conditions are taken from Skinner et al. (2019). The gravitational acceleration is constant with $g_{z}=-2$ and the initial density of the fluid is stratified such that:
$\rho(z)=\rho_{0}\,\Big{(}1-\frac{\gamma-1}{\gamma}\frac{\rho_{0}gz}{P_{0}}\Big{)}^{\frac{1}{\gamma-1}},\quad\rho_{0}=\left\{\begin{aligned} &\rho_{h}=3\quad\textrm{if}\quad z\geq 0,\\ &\rho_{l}=1\quad\textrm{otherwise.}\end{aligned}\right.$ (21)
@@
Figure 4: Multiepoch vertical profile of the horizontally-averaged passive scalar $\langle f_{\textsf{x}}\rangle$ obtained in the 3-D Rayleigh-Taylor instability test (see also Fig. 3, and §4.2 for discussion).
At $t=0$, the interface between the two fluids, which is originally at $z=0$, is displaced by an amount h(x,y) given by:
$\begin{split}h(x,y)&=\frac{1}{H}\sum_{8\leq k_{x}^{2}+k_{y}^{2}\leq 16}\Big{[}a\cos(k_{x}X)\cos(k_{y}Y)+b\cos(k_{x}X)\sin(k_{y}Y)\\ &+c\sin(k_{x}X)\cos(k_{y}Y)+d\sin(k_{x}X)\sin(k_{y}Y)\Big{]}\end{split}$ (23)
where $X=2\pi x/L$, $Y=2\pi y/L$ and $k_{x},k_{y}$ are the wave numbers with values in the range zero to four. The coefficients $a$, $b$, $c$ and $d$ are sampled from a uniform distribution randomly sampled from -1 to 1, excluding the boundaries. The normalized coefficient is $H=\sqrt{(1/4)(a^{2}+b^{2}+c^{2}+d^{2})}/(h_{\rm rms})$ and $h_{\rm rms}=3\times 10^{-4}L$.
@@
We can quantify the mixing between the two fluids by evaluating the horizontally-averaged concentration of the passive scalar versus $z$ defined as:
@@
The vertical profile of the quantity $\langle f_{\textsf{x}}\rangle$ is shown at multiple epochs in Fig. 4. Initially, the interface between the two fluids at $z\sim 0$ is sharp. However, as time passes, the mixing layer thickens and eventually encompasses the full height of the simulated domain. This mixing occurs on large scale and corresponds to advection of material with distinct $f_{\textsf{x}}$, whose average value covers from 0 to 1 over the full height.

### v2

@@ 2.5 Monitoring of internal energy in highly supersonic regions
The evolution of the internal energy is then given by
@@
The condition for $\alpha$ is
$\alpha=\left\{\begin{aligned} &1\quad&\textrm{if}\quad e\geq\varepsilon\,e_{k}\\ &\Big{(}\frac{e}{\varepsilon\,e_{k}}\Big{)}^{2}\quad&\textrm{if}\quad\varepsilon\,e_{k}>e>\frac{\varepsilon\,e_{k}}{10}\\ &0\quad&\textrm{if}\quad e\leq\frac{\varepsilon\,e_{k}}{10}\end{aligned}\right.,$ (12)
@@ 3.1 Numerical discretization
Our notation for space-time discretization is such that the position at the center of the cell is denoted by $r_{i}$ and the mesh nodes are given by $r_{i\pm 1/2}$. The cell size is $\delta r$. The time interval between $t^{n}$ and $t^{n+1}$ is $\delta t$. The notation $U^{n}_{i}$ represents the average quantity at time $t^{n}$ for the cell centered at $r_{i}$. This notation is trivially extended to all geometries and in several dimensions. When using spherical coordinates, $i$ refers to the radius coordinate, $j$ to the polar-angle coordinate, and $k$ to the azimuthal-angle coordinate.
@@ 3.1.1 Mesh
In HERACLES++, the computational domain is defined as a logically Cartesian mesh that can be either in Cartesian or spherical coordinates and in 1D, 2D, or 3D. The surface elements ($dS$) and volumes ($dV$) of cells are computed based on the chosen geometry. The grid spacing in each direction can be either regular or logarithmic. Users also have the option to implement a custom grid spacing. It is also possible to initialize the simulation from an HDF5 file containing both the physical quantities and the grid spacing.
In addition, the code includes a moving grid feature that allows one to shift the grid along a specified coordinate axis during the simulation. This functionality allows one to track regions of interest defined by the user. In the context of SN ejecta, the grid can be allowed to move in the radial direction in order to follow the expanding ejecta. With this option, the grid size can be reduced or resolution can be enhanced in specific regions. Furthermore, the elimination of cells in the inner ejecta reduces the constraints on the CFL conditions.
@@ 3.1.2 Momentum equation discretization in spherical coordinates
Equations 1 are written in a conservative form. However, in spherical coordinates, the tensor operator (Eq. 1b) presents geometrical source terms. We use a classical discretization method that conserves angular momentum. The azimuthal component, $\phi$, of the momentum equation is then written as:
$\begin{split}\frac{\partial\rho u_{\phi}}{\delta t}+\frac{1}{r^{2}}\frac{\partial r^{2}\rho u_{\phi}u_{r}}{\partial r}+\frac{1}{r\sin\theta}\frac{\partial\sin\theta\,\rho u_{\phi}u_{\theta}}{\partial\theta}\\ +\frac{1}{r\sin\theta}\frac{\partial(\rho u_{\phi}^{2}+P_{\phi\phi})}{\partial\phi}=-\,\Big{(}\frac{\rho u_{\phi}u_{r}}{r}+\frac{\cot\theta\,\rho u_{\phi}u_{\theta}}{r}\Big{)}\end{split}.$ (13)
@@
[…]
where $dS^{r}$ is the surface elements in the radial direction, $dS^{\theta}$ in the polar direction, and $dV$ the volume element. The values at the cell interfaces are computed using outputs from the Riemann solver used for the cell update.
The corresponding equation for the polar component, $\theta$, is
$\begin{split}\frac{\partial\rho u_{\theta}}{\delta t}+\frac{1}{r^{2}}\frac{\partial r^{2}\rho u_{\theta}u_{r}}{\partial r}+\frac{1}{r\sin\theta}\frac{\partial\sin\theta(\rho u_{\theta}^{2}+P_{\theta\theta})}{\partial\theta}\\ +\frac{1}{r\sin\theta}\frac{\partial\rho u_{\theta}u_{\phi}}{\partial\phi}=-\,\Big{(}\frac{\rho u_{\theta}u_{r}}{r}-\cot\theta\,\frac{\rho u_{\phi}^{2}}{r}\Big{)}\end{split},$ (15)
@@
$\begin{split}\cot\theta\,\frac{\rho u_{\phi}^{2}}{r}=\frac{([\rho u_{\phi}^{2}]_{k+1/2}+[\rho u_{\phi}^{2}]_{k-1/2})\cos[(\theta_{j+1/2}+\theta_{j-1/2})/2]}{2\sin[(\theta_{j+1/2}+\theta_{j-1/2})/2]}\\ \times\frac{(dS_{j+1/2}^{\theta}-dS_{j-1/2}^{\theta})}{2\,dV}\end{split}.$ (16)
@@ 3.2 Algorithmic steps
HERACLES++ was designed to easily incorporate a variety of numerical schemes. In this paper, we concentrate on a Godunov-type approach (Godunov & Bohachevsky, 1959) and more specifically on the Monotone Upstream-centered Schemes for Conservation Laws (MUSCL)-Hancock method (van Leer, 1977a, b, 1979, 1984; Toro, 2009) to solve the hydrodynamical equations. This method is a simple but practical choice for achieving second-order accuracy in numerical simulations. Its ability to accurately represent gradients and discontinuities, its robustness, and its conservation properties make it a reliable tool to treat complex flows involving strong shocks.
@@
$\tilde{U}^{n}_{L}=\tilde{U}^{n}_{i}-\frac{\delta r_{i}}{2}\sigma\quad\textrm{and}\quad\tilde{U}^{n}_{R}=\tilde{U}^{n}_{i}+\frac{\delta r_{i}}{2}\sigma,$ (17)
where $\tilde{U}_{L}$ and $\tilde{U}_{R}$ are the reconstructed primitive variables ($\rho,\vec{u},P$) at the left and right sides of the cell, and $\sigma$ the slope defined using a slope limiter. Three different slope limiters are implemented in HERACLES++: Minmod (Sandham & Yee, 1989), Van Leer, and Van Albada (Berger & Aftosmis, 2005; Toro, 2009). The user can choose the slope limiter at runtime and it is also possible to remain at first order. From these primitive reconstructed variables, we can obtain the reconstructed conservative variables, $U_{L,R}$.
@@
$U^{n+1/2}_{L,R}=U^{n}_{L,R}+\frac{\delta t}{2\delta r_{i}}\Big{[}F(U^{n}_{L})-F(U^{n}_{R})\Big{]},$ (18)
where $F$ is the flux. The fluxes are computed from these space-time extrapolated values using the analytical flux function. In 3D, the reconstruction is done in a similar manner (e.g., see Sect. 16.5 of Toro 2009).
@@
$U_{i}^{n+1}=U_{i}^{n}+\frac{\delta t}{\delta r}\Big{(}\mathcal{F}_{i-1/2}^{n+1/2}-\mathcal{F}_{i+1/2}^{n+1/2}\Big{)},$ (19)
@@ 3.3 Implementation
HERACLES++  is designed with an emphasis on both performance and portability. This enables the use of a single source code that can be compiled across various architectures (GPUs, CPUs), while ensuring optimal performance on each. The code is modular, making it easy to implement and switch between different slope limiters and Riemann solvers for the flux calculation.
@@ 3.3.1 Parallelism
The code can leverage MPI for distributed memory parallelism and the Kokkos library (Trott et al., 2022) for shared memory parallelism. The global computational domain is partitioned into equal subdomains, with each subdomain assigned to a separate MPI process. Within each subdomain, the Kokkos library is used to take advantage of shared memory parallelism.
@@
where U_new, U, and S are all of the Kokkos::View type, which is the fundamental data type used in the Kokkos library. The MDRangePolicy specifies how to iterate over a multidimensional container.
@@ 3.3.2 Inputs and outputs
The input-output (I/O) operations in HERACLES++ are handled by PDI, [^2] PDI stands for Portable Data Interface. Information can be found at https://pdi.dev/master/ a library designed to decouple high-performance simulation codes from I/O management. The simulation generates a significant amount of data, which is output by PDI into HDF5 files. We can initialize a simulation by directly assigning values to the physical fields or by reading an HDF5 file containing the values for these fields. This approach allows us to restart a simulation from a specific checkpoint by loading the corresponding data file.
@@ 4 Benchmarking tests
We have conducted a large number of tests to validate and benchmark HERACLES++. These include the standard suite of hydrodynamical tests for shocks, advection, implosion, or explosion. In Appendix C, we present results for a subset of such tests, in particular those that were performed in 2D and 3D. In this section, we present results for the tests that have a more direct relevance to the context of multidimensional simulations of stellar explosions. Hence, we present results for a 2D Kelvin-Helmholtz instability test in Sect. 4.1, for a 3D Rayleigh-Taylor instability test in Sect. 4.2, and for a 3D Sedov blast wave test in Sect. 4.3. We then move on to discuss a more astrophysics-oriented test in Sect. 5 with the simulations in 1D and 3D of the propagation of a SN shock in a RSG star envelope.
Figure 1: Color map of the density for the 2D Kelvin-Helmholtz instability test and shown at times 00 (left), $3$ (middle) and $6$ (right). The grid resolution corresponds to $[n_{x},n_{z}]=[2048,1024]$. (See Sect. 4.1 for discussion.)
@@ 4.1 Two-dimensional Kelvin-Helmholtz instability test
The Kelvin-Helmholtz instability test is useful to investigate the development of turbulence arising from a velocity shear within a fluid. Following Lecoanet et al. (2016), the initial conditions in the $(x,z)$ plane for the density, $\rho$, the $x$-velocity, $u_{x}$, and the $z$-velocity, $u_{z}$, are given by
@@
with $a=0.05$, $\sigma=0.2$, $z_{1}=0.5$, and $z_{2}=1.5$. The initial pressure, $P_{0}$, is 10 and we assume an ideal gas with $\gamma=5/3$. The boundary conditions are periodic in every direction.
Figure 2: Instability growth rate for the 2D Kelvin-Helmholtz instability test shown in Fig. 1. This growth rate is estimated from the evolution of the quantity $\partial\ln u_{z}^{max}/\partial t$ and for three different grid choices, corresponding to “Low” (blue), “Med” (red), and “High” (green) resolution (i.e., $[n_{x},n_{z}]=[512,256]$, $[1024,512]$, and $[2048,1024]$). The dashed line corresponds to the best-fitting exponential to the growth rate in the linear regime, which breaks down at $t\sim 2.6$.
Figure 3: Passive scalar concentration, $f_{\textsf{x}}$, for the 3D Rayleigh-Taylor instability test at a time of $5$ (left) and $10$ (right). The Cartesian grid is characterized by $[n_{x},n_{y},n_{z}]=[512,512,1024]$. (See Sect. 4.2 for discussion.)
Figure 1 shows the evolution of the density at three times ($t=0,3,6$). The instability results in the creation of small, alternate vortices, which cascade into turbulence. In Fig. 2, we quantify the growth rate of the instability-generated structures by evaluating the evolution of the quantity $\partial\ln u_{z}^{max}/\partial t$, with $u_{z}^{max}$ the maximum velocity at each time, and for different resolutions: $512\times 256$ (blue), $1024\times 512$ (red), and $2048\times 1024$ (green). The linear phase of the instability ends at $t\sim 2.6$ and the dashed black line represents the best exponential fitting for this phase, corresponding to $0.01e^{1.9t}$. The growth rate flattens after that time and remains roughly constant until the end of the simulation, largely independent of the resolution.
We can compare our results to those published by Skinner et al. (2019). The linear evolution that they obtain ends around a time of 2.5 with a value of 1.5, which are similar to our results. Their growth rate is between 1.5 and 2, whereas ours is a little lower between 1.2 and 1.7. When comparing Fig. 1 with the counterpart in their work, we observe fewer structures and a more diffusive scheme, which can explain the lower values in our growth rate. In both simulations, the resolution modestly affects the development of turbulence. Indeed, we only obtain a slight increase in the growth rate when employing a higher resolution.
@@ 4.2 Three-dimensional Rayleigh-Taylor instability test
The first three-dimensional test that we present is the development of the nonlinear R-T instability. The initial conditions are taken from Skinner et al. (2019). The gravitational acceleration is constant with $g_{z}=-2$ and the initial density of the fluid is stratified such that:
$\rho(z)=\rho_{0}\,\Big{(}1-\frac{\gamma-1}{\gamma}\frac{\rho_{0}gz}{P_{0}}\Big{)}^{\frac{1}{\gamma-1}},\quad\rho_{0}=\left\{\begin{aligned} &\rho_{h}=3\quad\textrm{if}\quad z\geq 0\\ &\rho_{l}=1\quad\textrm{otherwise}\end{aligned}\right.,$ (21)
@@
Figure 4: Multiepoch vertical profile of the horizontally averaged passive scalar $\langle f_{\textsf{x}}\rangle$ obtained in the 3D Rayleigh-Taylor instability test (see also Fig. 3, and Sect.4.2 for discussion).
At $t=0$, the interface between the two fluids, which is originally at $z=0$, is displaced by an amount, h(x,y), given by
$\begin{split}h(x,y)&=\frac{1}{H}\sum_{8\leq k_{x}^{2}+k_{y}^{2}\leq 16}\Big{[}a\cos(k_{x}X)\cos(k_{y}Y)+b\cos(k_{x}X)\sin(k_{y}Y)\\ &+c\sin(k_{x}X)\cos(k_{y}Y)+d\sin(k_{x}X)\sin(k_{y}Y)\Big{]}\end{split},$ (23)
where $X=2\pi x/L$, $Y=2\pi y/L$ and $k_{x},k_{y}$ are the wave numbers with values in the range zero to four. The coefficients, $a$, $b$, $c$, and $d$, are sampled from a uniform distribution randomly sampled from -1 to 1, excluding the boundaries. The normalized coefficient is $H=\sqrt{(1/4)(a^{2}+b^{2}+c^{2}+d^{2})}/(h_{\rm rms})$ and $h_{\rm rms}=3\times 10^{-4}L$.
@@
We can quantify the mixing between the two fluids by evaluating the horizontally averaged concentration of the passive scalar versus $z$ defined as:
@@
The vertical profile of the quantity, $\langle f_{\textsf{x}}\rangle$, is shown at multiple epochs in Fig. 4. Initially, the interface between the two fluids at $z\sim 0$ is sharp. However, as time passes, the mixing layer thickens and eventually encompasses the full height of the simulated domain. This mixing occurs on large scale and corresponds to advection of material with distinct $f_{\textsf{x}}$, whose average value covers from 0 to 1 over the full height.